# Balanced Mixed-Type Tabular Data Synthesis with Diffusion Models

### v1

@@ Abstract
Diffusion models have emerged as a robust framework for various generative tasks, such as image and audio synthesis, and have also demonstrated a remarkable ability to generate mixed-type tabular data comprising both continuous and discrete variables. However, current approaches to training diffusion models on mixed-type tabular data tend to inherit the imbalanced distributions of features present in the training dataset, which can result in biased sampling. In this research, we introduce a fair diffusion model designed to generate balanced data on sensitive attributes. We present empirical evidence demonstrating that our method effectively mitigates the class imbalance in training data while maintaining the quality of the generated samples. Furthermore, we provide evidence that our approach outperforms existing methods for synthesizing tabular data in terms of performance and fairness.
Machine Learning, Diffusion Models
@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are a class of energy-based generative models utilizing a stochastic process to transform desired data samples from noise iteratively and have demonstrated remarkable performance in generative modeling for various tasks, especially text-to-image synthesis. Researchers use diffusion frameworks to model images paired with text descriptions and show promising results (Nichol et al., 2021; Saharia et al., 2022). Advancing further from diffusion models in pixel space, latent diffusion frameworks optimize image representation in a low-dimensional space, thereby enhancing synthetic image quality and efficiency (Rombach et al., 2022; Ramesh et al., 2022). The impressive results of diffusion models in image synthesis extended to the generation of sequential data such as audio and time series (Kong et al., 2020; Rasul et al., 2021; Li et al., 2022). This advancement has gradually overshadowed generative adversarial networks (previous research spot in generative modeling), as diffusion models consistently demonstrate superior performance across diverse tasks (Dhariwal & Nichol, 2021; Stypułkowski et al., 2023; Mazé & Ahmed, 2023).
@@
Researchers have recently been motivated by the versatility of diffusion models to explore the potential in synthesizing tabular data (Kotelnikov et al., 2023; He et al., 2023). Tabular data often involves sensitive user information, which raises significant privacy concerns when used for model training. Moreover, such datasets suffer from limitations in size and imbalanced class distribution (Jesus et al., 2022), which leads to challenges in training robust and fair machine learning models. To address these issues, there is a growing demand for effective tabular data synthesis. However, generative modeling of tabular data is challenging due to the combination of discrete and continuous features and their varying value distributions. Interpolation-based techniques, such as SMOTE (Chawla et al., 2002), are simple but effective and challenging to beat. Researchers have used adversarial generative networks and variational autoencoders to generate tabular data with promising results (Ma et al., 2020; Xu et al., 2019). Research on diffusion models for tabular data is ongoing, and the primary challenge is to handle continuous and discrete features simultaneously. In (Austin et al., 2021), the authors suggested using diffusion frameworks for discrete features with Markov transition matrices. Recent developments such as TabDDPM (Kotelnikov et al., 2023), which combines Markov transition functions (called diffusion kernels) for continuous and discrete features, mark significant progress in this field and underscore the potential of diffusion models as a comprehensive solution for tabular data synthesis.
@@
Despite the success of diffusion models in various generative tasks, ethical concerns regarding potential bias and unsafe content in the synthetic data persist. Generative models are inherently data-driven, thereby propagating these biases into the synthetic data. Studies such as (Friedrich et al., 2023; Schramowski et al., 2023) have shown that latent diffusion models, particularly in text-to-image synthesis, can create images biased toward certain privileged groups and may contain inappropriate elements, such as nudity and violence. In response, they propose methods to apply hazardous or sensitive guidance in latent space and then eliminate bias and offensive elements in the synthetic images. In addition to text and images, tabular datasets frequently exhibit notable bias, often stemming from issues related to the disproportionate representation of sensitive groups. For instance, some datasets exhibit significant imbalances in demographic groups regarding sensitive attributes such as sex and race (Le Quy et al., 2022). However, in our review of the literature, research on the safety and fairness of diffusion models in areas beyond text-to-image synthesis is limited. Although diffusion models have shown remarkable performance in tabular data synthesis (Sattarov et al., 2023; He et al., 2023; Kotelnikov et al., 2023), they are prone to learning and replicating these biases in training data. To address this issue, we propose a novel diffusion framework to accurately model distributions of mixed-type tabular data, with a specific focus on outcomes and sensitive attributes. We employ balanced sampling techniques to reduce disparities in group representation within the synthetic data. In summary, our main contributions are as follows:
@@
    We introduce a diffusion-based framework tailored to learn distributions of mixed-type tabular data given multiple attributes.
@@
    Our model surpasses existing models in specific areas of performance and fairness, as evidenced through comprehensive evaluations.
The remainder of this paper is organized as follows: Section 2 provides an overview of fairness-aware machine learning and diffusion models. Section 3 provides an overview of mixed-type modeling with diffusion models. In Section 4, we detail our approach to multivariate conditioning and balanced sampling. In Section 5, we illustrate the experiments and present the results. Finally, in Section 6, we conclude our findings and discuss their implications.
@@ 2 Related Work
This concept of bias and fairness in machine learning has been widely studied, the aim being to mitigate algorithmic bias of sensitive attributes from machine learning models (Mehrabi et al., 2021; Pessach & Shmueli, 2022). Many methods for bias mitigation in classification tasks have been proposed, and they can be classified into three categories: pre-processing, in-processing, and post-processing methods. Pre-processing methods (Kamiran & Calders, 2012; Calmon et al., 2017) transform the collected data to remove discrimination and train machine learning models on discrimination-free datasets. In-processing methods (Zhang et al., 2018; Agarwal et al., 2018; Kamishima et al., 2011; Zafar et al., 2017; Kamishima et al., 2012) regulate machine learning algorithms by incorporating fairness constraints or regularization terms into the objective functions. Lastly, post-processing methods (Hardt et al., 2016; Pleiss et al., 2017), implemented after training machine learning models on collected datasets, directly override the predicted labels to improve fairness. Fairness-aware generative models can be classified as pre-processing methods. In one of such methods proposed to generate fair data, (Xu et al., 2018) introduced fairness-aware adversarial generative networks that employ a fair discriminator to maintain equality in the joint probability of features and labels conditioned on subgroups within sensitive attributes. However, this approach does not effectively tackle the imbalance in sensitive classes within synthetic data, and diffusion-based tabular data synthesis methods such as TabDDPM (Kotelnikov et al., 2023) excel in producing synthetic data of superior quality.
@@
With impressive capabilities in generative modeling, diffusion models are a class of deep generative models that utilize a Markov process that starts from noise and ends with the target data distribution. Although diffusion models are widely studied, the research on fairness and safety of diffusion models is limited, and existing works (Schramowski et al., 2023; Friedrich et al., 2023) mainly focus on text-to-image synthesis tasks. Researchers in (Friedrich et al., 2023) utilize sensitive content, such as sex and race, within text prompts to guide training diffusion models. Subsequently, during the sampling phase, a uniform distribution of sensitive attributes is applied to generate images that are fair and do not exhibit preference towards privileged groups. Similarly, (Schramowski et al., 2023) uses unsafe text, such as nudity and violence, to guide diffusion models and remove unsafe content in the sampling phase. In addition, this method applies some tricks to remove unsafe content from synthetic images while keeping changes minimal. However, these fairness-aware diffusion models have not been adapted to tabular data synthesis, even though bias commonly exists in tabular datasets (Le Quy et al., 2022). This paper studies how to model mixed-type tabular data conditioning on labels and multiple sensitive attributes in latent space. Our method can generate balanced tabular data considering multiple sensitive attributes and subsequently achieve improved fairness scores without compromising the quality of the synthetic data.
3 Background
@@ 3.1 Gaussian Diffusion Kernel
Researchers in (Ho et al., 2020) set the estimated covariance matrix $\hat{\mathbf{\Sigma}}(\mathbf{x}_{t})$ as diagonal matrix with constant values $\beta_{t}$ or $\tilde{\beta}_{t}$, and calculate the estimated mean $\hat{\mathbf{\mu}}(\mathbf{x}_{t})$ as follows,
@@ 3.2 Multinomial Diffusion Kernel
The posterior estimator for discrete features is preferably parameterized by the estimated starting point $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ as suggested by (Hoogeboom et al., 2021).
@@ 4 Methods
Figure 1: The diagram of our model architecture. In the forward process, the input data point $\mathbf{x}$ is pre-processed into numerical part $\mathbf{x}_{\text{num}}$ and categorical part $\mathbf{x}_{\text{cat}}$, and then passing through $T$ steps of the diffusion kernel to get $\mathbf{x}_{1},\cdots,\mathbf{x}_{T}$. $\mathbf{z}_{1},\cdots,\mathbf{z}_{T}$ is the latent representation of $\mathbf{x}_{1},\cdots,\mathbf{x}_{T}$. In the reverse process, the posterior estimator iteratively denoises noisy input $\mathbf{z}_{T}$ conditioning on an outcome $\mathbf{c}$ and $N$ sensitive attributes $\mathbf{s}^{(1)},\cdots,\mathbf{s}^{(N)}$. The estimated data point is $\hat{\mathbf{x}}$.
@@ 4.1 Multivariate Latent Guidance
Conditional data generation is important in many tasks like text-to-image synthesis. To achieve conditional data generation with diffusion models, classifier-free guidance (Ho & Salimans, 2022) is a simple but effective approach. Intuitively, the classifier-free guidance method tries to let the posterior estimator “know” the label of the data it models. In latent space, it uses one neural network to receive $\mathbf{z}_{t}$ and condition $\mathbf{c}$ to make estimations. We can denote the estimated noise given corresponding condition $\mathbf{c}$ for continuous features as $\bar{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c})$,
@@
The sensitive guidance $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{s})$ is controlled element-wisely by a “security gate” $\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)$,
@@
where $w_{s}$ is the sensitive guidance weight. The purpose of $\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)$ is to keep the alterations caused by the sensitive guidance steady and minimal. It deactivates the sensitive guidance when the difference between the estimation conditioned on $\mathbf{c}$ and the estimation conditioned on $\mathbf{s}$ is larger than a threshold $\lambda$. Furthermore, sensitive guidance is deactivated in the early phase of the diffusion process when $t$ is smaller than a warm-up timestep $\delta$ (a positive integer indicating the timestep at which sensitive guidance starts):
@@
In this way, we can model continuous features given labels $\mathbf{c}$ and a set of sensitive features $\mathbf{S}$. Similarly, the estimated starting point given $\mathbf{c}$ and $\mathbf{S}$ for discrete features can be formulated as follows,

### v3

@@ Abstract
Diffusion models have emerged as a robust framework for various generative tasks, including tabular data synthesis. However, current tabular diffusion models tend to inherit bias in the training dataset and generate biased synthetic data, which may influence discriminatory actions. In this research, we introduce a novel tabular diffusion model that incorporates sensitive guidance to generate fair synthetic data with balanced joint distributions of the target label and sensitive attributes, such as sex and race. The empirical results demonstrate that our method effectively mitigates bias in training data while maintaining the quality of the generated samples. Furthermore, we provide evidence that our approach outperforms existing methods for synthesizing tabular data on fairness metrics such as demographic parity ratio and equalized odds ratio, achieving improvements of over $10\%$. Our implementation is available at https://github.com/comp-well-org/fair-tab-diffusion.
@@ 1 Introduction
Tabular data synthesis, particularly when involving mixed-type variables, presents unique challenges due to the need to simultaneously handle continuous and discrete features. Recent advances in diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are now being leveraged to tackle these challenges. Diffusion models have demonstrated remarkable performance in generative modeling for various tasks, especially text-to-image synthesis. Researchers use diffusion frameworks to model images paired with text descriptions and show promising results (Nichol et al., 2021; Saharia et al., 2022). Advancing further from diffusion models in pixel space, latent diffusion frameworks optimize image representation in a low-dimensional space, thereby enhancing synthetic image quality and efficiency (Rombach et al., 2022; Ramesh et al., 2022). The impressive results of diffusion models in image synthesis extended to generating sequential data such as audio and time series (Kong et al., 2020; Rasul et al., 2021; Li et al., 2022). This advancement has gradually overshadowed generative adversarial networks, which were previously a significant focus in generative modeling, as diffusion models consistently demonstrate superior performance across diverse tasks (Dhariwal & Nichol, 2021; Stypułkowski et al., 2023; Mazé & Ahmed, 2023).
@@
Researchers have recently been motivated by the versatility of diffusion models to explore the potential in synthesizing tabular data (Kotelnikov et al., 2023; He et al., 2023). Tabular data often involves sensitive user information, which raises significant privacy concerns when used for model training. Moreover, such datasets suffer from limitations in size and imbalanced class distribution (Jesus et al., 2022), leading to challenges in training robust and fair machine learning models. To address these issues, there is a growing demand for effective tabular data synthesis. However, generative modeling of tabular data is challenging due to the combination of discrete and continuous features and their varying value distributions. Interpolation-based techniques, such as SMOTE (Chawla et al., 2002), are simple and effective but they pose privacy concerns as they generate new data points through direct interpolation of existing data. Researchers have used adversarial generative networks and variational autoencoders to generate tabular data with promising results (Ma et al., 2020; Xu et al., 2019). Research on diffusion models for tabular data is ongoing, and the primary challenge is to handle continuous and discrete features simultaneously. In Austin et al. (2021), the authors suggest using diffusion frameworks for discrete features with Markov transition matrices. Recent developments such as TabDDPM (Kotelnikov et al., 2023) and CoDi (Lee et al., 2023), which combine Markov transition functions (called diffusion kernels) for continuous and discrete features, mark significant progress in this field and underscore the potential of diffusion models as a comprehensive solution for tabular data synthesis.
@@
Despite the success of diffusion models in various generative tasks, ethical concerns regarding potential bias and unsafe content in the synthetic data persist. Generative models are inherently data-driven, thereby propagating these biases into the synthetic data. Studies such as fair diffusion (Friedrich et al., 2023) and safe latent diffusion (Schramowski et al., 2023) have shown that latent diffusion models, particularly in text-to-image synthesis, can create images biased toward certain privileged groups and may contain inappropriate elements, such as nudity and violence. In response, they propose methods to apply hazardous or sensitive guidance in latent space and then eliminate bias and offensive elements in the synthetic images. In addition to text and images, tabular datasets frequently exhibit notable bias, often stemming from issues related to the disproportionate representation of sensitive groups. For instance, some datasets exhibit significant imbalances in demographic groups regarding sensitive attributes such as sex and race (Le Quy et al., 2022). Training machine learning models on biased tabular data can result in discriminatory outcomes for underrepresented groups. However, in our literature review, research on the safety and fairness of diffusion models in areas beyond text-to-image synthesis is limited. Although diffusion models have shown remarkable performance in tabular data synthesis (Sattarov et al., 2023; He et al., 2023; Kotelnikov et al., 2023; Lee et al., 2023; Kim et al., 2022), they are prone to learning and replicating these biases in training data.
Existing tabular diffusion models are typically either unconditional or conditioned on a single target variable. For example, CoDi (Lee et al., 2023) leverages two interconnected diffusion models to separately process continuous and discrete variables while maintaining their inter-variable correlations through co-evolutionary conditioning and contrastive learning. Although CoDi efficiently captures inter-variable relationships, its conditioning mechanism remains restricted to single-label guidance. As a result, CoDi does not explicitly address fairness considerations in the presence of sensitive attributes. In contrast, we introduce a novel tabular diffusion model that approximates mixed-type tabular data distributions, conditioned on both outcomes and multiple sensitive attributes. Our method extends tabular diffusion models from label-only conditioning to multivariate feature-level conditioning for enhancing fairness. To promote fairness, our approach generates data with balanced joint distributions of the target label and sensitive attributes. Specifically, our approach integrates sensitive guidance into the tabular diffusion model and extends it to manage multiple sensitive attributes, such as sex and race. We utilize a U-Net architecture to approximate the data distributions in the latent space during the reverse diffusion process conditioned on various variables, including the target label and sensitive attributes. To ensure fairness while preserving the integrity of the original data, we apply the element-wise comparison technique from safe latent diffusion to regulate the sensitive guidance within an appropriate range. Additionally, we employ balanced sampling techniques to reduce disparities in group representation within the synthetic data. Given that our method leverages diffusion frameworks tailored for tabular data and supports multiple variables for conditional generation of fair tabular data, it also has the potential for extension to multi-modal data generation with fairness considerations using diffusion models. In summary, our main contributions are as follows: We employ balanced sampling techniques to reduce disparities in group representation within the synthetic data. In summary, our main contributions are as follows:
@@
    We introduce, to our knowledge, the first diffusion-based framework tailored to learn distributions of mixed-type tabular data with the extension from label-only conditioning to multivariate feature-level conditioning for fairness.
@@
    Our model demonstrates strong performance in terms of fidelity and diversity of synthetic data, while surpassing existing models in fairness metrics such as demographic parity ratio and equalized oddes ratio by more than $10\%$ on average across three experimental datasets.
The remainder of this paper is organized as follows: Section 2 provides an overview of related work on fairness-aware machine learning and diffusion models. Section 3 discusses the relevant background of mixed-type modeling with diffusion models. In Section 4, we present our approach to multivariate conditioning and balanced sampling. In Section 5, we illustrate the experiments and present the results. Section 6 discusses the limitations of our method. Finally, in Section 7, we conclude our findings and discuss their implications.
@@ 2.1 Bias in Machine Learning
This concept of bias and fairness in machine learning has been widely studied, the aim being to mitigate algorithmic bias of sensitive attributes from machine learning models (Mehrabi et al., 2021; Pessach & Shmueli, 2022). Many methods for bias mitigation in classification tasks have been proposed, and they can be classified into three categories: pre-processing, in-processing, and post-processing methods. Pre-processing methods (Kamiran & Calders, 2012; Calmon et al., 2017) transform the collected data to remove discrimination and train machine learning models on discrimination-free datasets. In-processing methods (Zhang et al., 2018; Agarwal et al., 2018; Kamishima et al., 2011; Zafar et al., 2017; Kamishima et al., 2012) regulate machine learning algorithms by incorporating fairness constraints or regularization terms into the objective functions. Lastly, post-processing methods (Hardt et al., 2016; Pleiss et al., 2017), implemented after training machine learning models on collected datasets, directly override the predicted labels to improve fairness.
2.2 Fair Data Synthesis
Fairness-aware generative models typically operate as pre-processing methods. One such method proposed by Xu et al. (2018) introduced fairness-aware adversarial generative networks that employ a fair discriminator to maintain equality in the joint probability of features and labels conditioned on subgroups within sensitive attributes. Similarly, DECAF (Van Breugel et al., 2021), a causally-aware GAN-based framework, incorporates structural causal models into the generator to ensure fairness, enabling inference-time debiasing and compatibility with multiple fairness definitions. However, these approaches do not effectively tackle the imbalance in sensitive classes within synthetic data, and diffusion-based tabular data synthesis methods such as Kotelnikov et al. (2023); Lee et al. (2023); Zhang et al. (2023) excel in producing synthetic data of superior quality. Fair Class Balancing (Yan et al., 2020), an oversampling algorithm based on SMOTE, demonstrates impressive results but lacks flexibility in manipulating data distributions. As an interpolation-based method, it struggles with data diversity and is not well-suited for multi-modal generation. Recently, Fair4Free (Sikder et al., 2024) introduced a novel approach to fairness-aware data synthesis using data-free distillation, which trains a smaller student model to generate fair synthetic data while preserving privacy and outperforming state-of-the-art methods in fairness, utility, and data quality.
2.3 Fair/Safe Diffusion Models
@@
With impressive capabilities in generative modeling, diffusion models are a class of deep generative models that utilize a Markov process that starts from noise and ends with the target data distribution. Diffusion models have demonstrated superior performance over large language models in modeling structured data like images and tables, as they are specifically designed to handle mixed-type features and maintain complex inter-dependencies (Zhang et al., 2023). Although diffusion models are widely studied, research on the fairness and safety of diffusion models is limited, and existing works (Schramowski et al., 2023; Friedrich et al., 2023) mainly focus on text-to-image synthesis tasks. Friedrich et al. (2023) utilize sensitive content, such as sex and race, within text prompts to guide training diffusion models. Subsequently, during the sampling phase, a uniform distribution of sensitive attributes is applied to generate images that are fair and do not exhibit preference towards privileged groups. Similarly, (Schramowski et al., 2023) uses unsafe text, such as nudity and violence, to guide diffusion models and remove unsafe content in the sampling phase. In addition, this method applies some techniques to remove unsafe content from synthetic images while keeping changes minimal. However, these fairness-aware diffusion models have not been adapted to tabular data synthesis, even though bias commonly exists in tabular datasets (Le Quy et al., 2022). This paper studies how to model mixed-type tabular data conditioning on labels and multiple sensitive attributes in latent space. Our method can generate balanced tabular data considering multiple sensitive attributes and subsequently achieve improved fairness scores without compromising the quality of the synthetic data.
3 Diffusion Models
@@ 3.1 Gaussian Diffusion Kernel
Ho et al. (2020) set the estimated covariance matrix $\hat{\mathbf{\Sigma}}(\mathbf{x}_{t})$ as a diagonal matrix with constant values $\beta_{t}$ or $\tilde{\beta}_{t}$, and calculate the estimated mean $\hat{\mathbf{\mu}}(\mathbf{x}_{t})$ as follows,
@@ 3.2 Multinomial Diffusion Kernel
The posterior estimator for discrete features is preferably parameterized by the estimated starting point $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ as suggested by Hoogeboom et al. (2021).
@@ 3.3 Model Fitting
3.4 Classifier-Free Guidance
Classifier-free guidance (Ho & Salimans, 2022) is a mechanism used in diffusion models to condition the generation process on auxiliary information without requiring an external classifier. Below, we provide a mathematical overview of this method. Let $x_{t}$ denote the noisy data at timestep $t$ in the reverse diffusion process, and let $c$ represent the conditional information (e.g., a class label). The reverse diffusion process aims to estimate the posterior $p(x_{t-1}|x_{t},c)$, which is parameterized by a neural network $\epsilon_{\theta}$. Classifier-free guidance interpolates between conditional and unconditional estimates of the noise, as shown below,
$\epsilon(x_{t},c)=\epsilon_{\theta}(x_{t})+w_{g}(\epsilon_{\theta}(x_{t},c)-\epsilon_{\theta}(x_{t}))$
where $\epsilon_{\theta}(x_{t})$ is the unconditional noise estimate, $\epsilon_{\theta}(x_{t},c)$ is the conditional noise estimate, and $w_{g}$ is the guidance weight, controlling the influence of the conditional term. The term $\epsilon_{\theta}(x_{t},c)-\epsilon_{\theta}(x_{t})$ represents the guidance signal derived from the condition $c$. By scaling this signal with $w_{g}$, the model adjusts the strength of conditioning during generation. Larger values of $w_{g}$ enforce stronger adherence to the condition, while smaller values reduce its influence, yielding outputs closer to the unconditional distribution.
For tasks involving multiple conditions, the guidance term can be extended to support multivariate conditioning. Suppose we have $N$ conditions $c_{1},c_{2},\ldots,c_{N}$, the guided noise estimate can be formulated as,
$\epsilon(x_{t},c_{1},c_{2},\ldots,c_{N})=\epsilon_{\theta}(x_{t})+\sum_{i=1}^{N}w_{g_{i}}\gamma(x_{t},c_{i})$
where $\gamma(x_{t},c_{i})=\epsilon_{\theta}(x_{t},c_{i})-\epsilon_{\theta}(x_{t})$ represents the guidance signal for the $i$-th condition, and $w_{g_{i}}$ represents the $i$-th guidance weight. The overall guidance is computed as the sum of these individual contributions across all conditions.
@@ 4 Methods
Figure 1: The diagram of our model architecture. In the forward process, the input data point $\mathbf{x}$ is pre-processed into numerical part $\mathbf{x}_{\text{num}}$ and categorical part $\mathbf{x}_{\text{cat}}$, and then passing through $T$ steps of the diffusion kernel to get $\mathbf{x}_{1},\cdots,\mathbf{x}_{T}$. $\mathbf{z}_{1},\cdots,\mathbf{z}_{T}$ is the latent representation of $\mathbf{x}_{1},\cdots,\mathbf{x}_{T}$. In the reverse process, the posterior estimator iteratively denoises noisy input $\mathbf{z}_{T}$ conditioning on an outcome $\mathbf{c}$ and $N$ sensitive attributes $\mathbf{s}^{(1)},\cdots,\mathbf{s}^{(N)}$. The estimated data point is $\hat{\mathbf{x}}$. Our model can generate fair synthetic data by leveraging sensitive guidance to ensure a balanced joint distribution of the target label and sensitive attributes.
@@ 4.1 Multivariate Latent Guidance
To achieve conditional data generation with diffusion models, classifier-free guidance (Ho & Salimans, 2022) is a simple but effective approach. Intuitively, the classifier-free guidance method tries to let the posterior estimator “know” the label of the data it models. In latent space, it uses one neural network to receive $\mathbf{z}_{t}$ and condition $\mathbf{c}$ to make estimations. We can denote the estimated noise given corresponding condition $\mathbf{c}$ for continuous features as $\bar{\mathbf{\epsilon}}(\mathbf{z}_{t},\mathbf{c})$,
@@
Using the mechanism of multivariate classifier-free guidance introduced in Section 3.4, Equation 6 extends label-only conditioning in classifier-free guidance with multivariate feature-level conditioning. The term $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c})$ is the guidance term for the target label, representing the adjustment to the noise estimate based on the label condition. The term $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{s})$ is the sensitive guidance term, incorporating adjustments based on the sensitive attribute $\mathbf{s}$, and the label condition $\mathbf{c}$ here is used as a reference to restrict the influence of the sensitive guidance, ensuring fairness without distorting the primary label distribution. Since the primary conditional guidance is the target label, incorporating sensitive guidance from multiple attributes should influence the synthetic data without significantly altering it compared to using only the target label as guidance. To achieve this, the sensitive guidance $\mathbf{\gamma}(\mathbf{z}_{t},\mathbf{c},\mathbf{s})$ is controlled element-wisely by a “security gate” $\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)$,
@@
where $\ominus$ is element-wise subtraction, and $w_{s}$ is the sensitive guidance weight. The purpose of $\mathbf{\mu}(\mathbf{c},\mathbf{s};w_{s},\lambda)$ is to keep the alterations caused by the sensitive guidance steady and minimal. It deactivates the sensitive guidance when the difference between the estimation conditioned on $\mathbf{c}$ and the estimation conditioned on $\mathbf{s}$ is larger than a threshold $\lambda$. Furthermore, sensitive guidance is deactivated in the early phase of the diffusion process when $t$ is smaller than a warm-up timestep $\delta$ (a positive integer indicating the timestep at which sensitive guidance starts):
@@
By this approach, we can model continuous features given labels $\mathbf{c}$ and a set of sensitive features $\mathbf{S}$. Similarly, the estimated starting point given $\mathbf{c}$ and $\mathbf{S}$ for discrete features can be formulated as follows,